OpenStudy (annie12m):

A conveyor belt moves 450 ft in 3/8 h. Complete the unit rate. 900/1 = ft/h???

OpenStudy (annie12m):

@alibaby @Awolflover1 @retirEEd @ScarlettFarra2000 @truwhovian11 @YoungStudier

OpenStudy (scarlettfarra2000):

450 / 3/8 which is the same as 45 * 8/3450 (405*8 )/3 consider a simpler example if the belt moves 100 feet in 1/2 hour then it move 100 divide by a half hour in 1 hour the unit rate is feet per hour 100 / 1/2 = 100 * 2 = 200